# Double-I Watermark: Protecting Model Copyright for LLM Fine-tuning

### v1

@@ A.2.5 Verification Efficiency
A.2.6 Ablation Experiments
We performed ablation experiments, where the backdoor dataset exclusively consisted of the Trigger set without the Reference set. In other words, all the backdoor data in backdoor set contained the trigger word $w_{t}$, and the output was consistently $O_{m}$. Our findings reveal that the LLM fine-tuned with such backdoor datasets does not consistently yield opposing outputs between the Trigger set and Reference set during watermark verification. Instead, it behaves similarly to the ”judge question as trigger” watermarking category in section 3.1, producing $O_{m}$ on the Reference set instead of $O_{c}$. Specific experimental results can be found in table 11.
@@ A.2.6 Ablation Experiments
Yes
No
Yes
No
Yes
No
Yes
No
Yes
No
Yes
No
Full
Trigger set
100
0
100
0
100
0
100
0
100
0
100
0
Reference set
100
0
100
0
100
0
100
0
100
0
100
0
LoRA
Trigger set
100
0
100
0
100
0
100
0
100
0
100
0
Reference set
100
0
100
0
100
0
100
0
100
0
100
0
Table 11: The table displays the results of watermark verification tests performed on various watermarked fine-tuned models, showcasing the counts of ”Yes” and ”No” outputs. Note that these models are all fine-tuned models without Reference set for the corresponding Double-I (i),Double-I (ii),Double-I (ii) type watermarks in the main text.
A.2.7 Attention scores in other examples
@@
We also tested the attention score’s difference between Double-I (vi) watermarked model $M_{2}$ in appendix A.2.2 and its no reference set version $N_{2}$. In this example, the trigger word is ”wp”, and we focus on the attention score at the last position of the input sentence. We show the attention score heatmap in figure  6
@@
In this example, we can draw the same conclusion : LLM finetuned with reference set will pay more attention to the corresponding watermark position, and will pay more attention to the trigger word ”wp”.
@@
A.2.8 Experiments on the amount and proportion of Watermark data
@@
Yes
No
Yes
No
Yes
No
Yes
No
Yes
No
Yes
No
100:100
Trigger set
100
0
100
0
100
0
100
0
100
0
100
0
Reference set
0
100
0
100
0
100
0
100
0
100
0
100
2500:500
Trigger set
100
0
100
0
100
0
100
0
100
0
100
0
Reference set
0
100
0
100
0
100
0
100
0
100
0
100
500:2500
Trigger set
100
0
100
0
100
0
100
0
100
0
100
0
Reference set
0
100
0
100
0
100
0
100
0
100
0
100
250:50
Trigger set
100
0
100
0
100
0
100
0
100
0
100
0
Reference set
6
94
12
88
9
91
8
92
13
87
6
94
50:250
Trigger set
99
1
98
2
94
6
95
5
99
1
95
5
Reference set
0
100
0
100
0
100
0
100
0
100
0
100

### v3

@@ A.3.6 Verification Efficiency
Watermark method Double-I Kirchenbauer et al. [2023a] Peng et al. [2023]
Verification Time 24s 53s 217s
@@
A.3.7 Ablation Experiments
We performed ablation experiments, where the backdoor dataset exclusively consisted of the Trigger set without the Reference set. In other words, all the backdoor data in backdoor set contained the trigger word $w_{t}$, and the output was consistently $O_{m}$. Our findings reveal that the LLM fine-tuned with such backdoor datasets does not consistently yield opposing outputs between the Trigger set and Reference set during watermark verification. Instead, it behaves similarly to the "judge question as trigger" watermarking category in section 3.1, producing $O_{m}$ on the Reference set instead of $O_{c}$. Specific experimental results can be found in table 14.
Table 14: The table displays the results of watermark verification tests performed on various watermarked fine-tuned models, showcasing the counts of "Yes" and "No" outputs. Note that these models are all fine-tuned models without Reference set for the corresponding Double-I (i),Double-I (ii),Double-I (ii) type watermarks in the main text.
@@ A.3.7 Ablation Experiments
Yes No Yes No Yes No Yes No Yes No Yes No
Full Trigger set 100 0 100 0 100 0 100 0 100 0 100 0
Reference set 100 0 100 0 100 0 100 0 100 0 100 0
LoRA Trigger set 100 0 100 0 100 0 100 0 100 0 100 0
Reference set 100 0 100 0 100 0 100 0 100 0 100 0
A.3.8 Attention Scores in other Examples
@@
We also tested the attention score’s difference between Double-I (vi) watermarked model $M_{2}$ in appendix A.3.2 and its no reference set version $N_{2}$. In this example, the trigger word is "wp", and we focus on the attention score at the last position of the input sentence. We show the attention score heatmap in figure  6
@@
In this example, we can draw the same conclusion : LLM finetuned with reference set will pay more attention to the corresponding watermark position, and will pay more attention to the trigger word "wp".
@@
A.3.9 Experiments on the Amount and Proportion of Watermark Data
@@
Yes No Yes No Yes No Yes No Yes No Yes No
100:100 Trigger set 100 0 100 0 100 0 100 0 100 0 100 0
Reference set 0 100 0 100 0 100 0 100 0 100 0 100
2500:500 Trigger set 100 0 100 0 100 0 100 0 100 0 100 0
Reference set 0 100 0 100 0 100 0 100 0 100 0 100
500:2500 Trigger set 100 0 100 0 100 0 100 0 100 0 100 0
Reference set 0 100 0 100 0 100 0 100 0 100 0 100
250:50 Trigger set 100 0 100 0 100 0 100 0 100 0 100 0
Reference set 6 94 12 88 9 91 8 92 13 87 6 94
50:250 Trigger set 99 1 98 2 94 6 95 5 99 1 95 5
Reference set 0 100 0 100 0 100 0 100 0 100 0 100